TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MirkoPojer on 2007-05-04 - 23:08

Temperatures, humidity and condensation status in Sector 7-8, Fri 4.

Position T [C] T_meteo [C] RH_meteo [%] T_portable [C] RH_portable [%] Condensation Dew Point [C] Air Speed [m/s] Time
DFBMC   20.4 48.1         0.3-0.8 17.30
    20.4 48.3 20.1 51.1       19.30
Q6_L8 (120A) 8.5     19.7 53.1 Low 9.9    
Q8_L8 (120A) 9     19.8 53.2 Low 10.0    
Q9_L8 (120A) 10     19.3 54.4 No 10.0 1.1-1.4  
Q10_L8 (120A) 9     19.0 54.2 Low 9.5   18.50
Q11_L8 (60A) 9     19.0 54.6 No 9.5    
Q12_L8 (60A) 9 18.5 52 18.6 54.5 Yes 9.2 1.0-1.3  
Q13_L8 (60A) 10     18.4 55.2 No 9.2    
Q14_L8 (60A) 9     18.5 55.1 Yes 9.3    
Q15_L8 (60A) 8.5     18.2 55.9 Yes 9.2   19.15

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-04 - 19:27

Short Circuit Test in UA27- Schedule:

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA - 11-05-07
8h run 4-8kA - 29-05-07
Interlock tests RQD&F - 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests ON GOING 03-05-07
600ANOEE powering tests - 08-05-07
600A-EE interlock tests ON GOING 03-05-07
600A-EE powering tests - 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-04 - 18:55

SCT in UJ33

PowerPoint GlobalAnalysis24hHR_UJ33.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x05x04x18x27/GlobalAnalysis24hHR_UJ33.ppt

  • During the 24hr heat run major problems were found (for a more detailed description see the attached file):
1 - Over-heating of the systems in UJ33 (central part) – TS-EL will re-routing the cables inside the UJ33.

2 - Over- heating of the cables in the tunnel side RZ33 – a vertical cable tray will be positioned parallel to the TS-CV cooling water supports. The tray will host one of the three cable layers that are now placed on the same tray.

  • The re-routing cable works will last approximately up to the third week of June. The integration/cost studies are on going and the possible co-activities interfering with the preparation for the cooldown of the A34 are being taken into account.
A meeting to up-date the "status of the art" is foreseen for Monday 14th. Details will be given.

  • During the inspection of the UJ33 P. Proudlock commented:

1 - ODH system must be foreseen in UJ33 - the DFBLC inside the UJ makes it indispensable

2 - Rectangular ducts (for the geometers) must be closed during the heat-run

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 18:35

Programme for Monday 07.05.2007

Programme for Monday 07.05.2007

DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL)

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Unlock of the quench heater power supplies (DQHDS) of Q5 and charge. AT/MEL
PC Configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
P2N HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

WHEN CONDITIONS IN ARC ARE MET
go on with the programme for the 60A converters (orbit correctors) and 120A (Q7 to Q10 correctors) as well as ElQAs in DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Friday 04.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the power converters for these circuits (before 9:00) AB/PO
Verification of the connections at the DFBMC side AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-04 - 16:35

RAT meeting 4/5/2007

PowerPoint Current_lead_operating_conditions_20070504.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/Current_lead_operating_conditions_20070504.ppt
Presentation by A. Ballarino - conditions for powering
PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 16:03

Cooldown Sector 78

Despite the long series of faults (network, cooling water, pumping group) experienced at the end of last week, the cryo team managed to bring down the average temperature of the magnets to 3 K during the week-end. The switch over to the new cryoplant took place on Monday but the connection of the new pumping group was delayed until Wednesday morning. Stabilization of temperatures and pressures as well as fine adjustments of the plants took place most of Wednesday until Thursday at 11.30 when the pumping group was set at 15 mbar. Unfortunately, at 11.30 an attempt to restart the UPS system on the surface caused a power cut which stopped all the PLCs. The cryo team is still recovering from that failure and we expect to be at nominal cryogenic condition in the arc tonight Friday May 4th.

Following a top management meeting which took place on Monday, a series of measures towards the clearing of the difficulties with the commissioning of the DFBs were taken: two different ways to avoid condensation on the chimneys are now in place: blowers on 6kA leads + heaters on the 120A leads on the DFBMC and dry-air bags around the chimneys of the DFBMA.

Although the powering front was very quiet during the week, significant progress was achieved with the preparation of the sequencer for the independently powered quadrupoles and the 120A correctors: we now have the sequencer tuned for the 60 A and 120 A correctors as well as the independently powered quadrupoles and D2.

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login